# Approaching the Quantum Speed Limit in Quantum Gates with Geometric Control

### v1

@@ III.3 Optimization in a 4-dimensional space
These examples suggest that the“plane condition” associated with $f_{\rm Plane}(t)$ plays a more critical role than the“phase condition” associated with $f_{\rm Phase}(t)$. When considering these functions, an optimization involving only $f_{\rm Plane}(t)$ yields better results than one involving only $f_{\rm Phase}(t)$. Nevertheless, as expected, the optimization procedure results in better trade-offs between QSL and quantum fidelity when both functions are minimized simultaneously. The performance of the procedure improves with the number of available control parameters, as the geometric optimization routine can then explore a broader range of quantum paths.
Figure 2: QSL efficiency after geometric optimization with a limited Hamiltonian control: (a) Driving without QSL optimization ($p_{1}=p_{2}=0$). (b): Optimization of the driving with only the plane condition ($p_{1}=0.5,p_{2}=0$). (c): Optimization of the driving with only the phase condition ($p_{1}=0$,$p_{2}=0.1$) (d): Optimization of the driving with both conditions ($p_{1}=0.5$,$p_{2}=0.1$). The 7-dimensional control vector, $\mathbf{v}(t)=(\Delta_{11}(t),\Omega_{ij}^{R}(t))$ with an alphabetic ordering for $(i,j)\in\{1,2,3,4\}^{2}$ and $i<j$ is initialized using $\mathbf{v}^{(0)}(t)=v_{0}(1,1,1,1,1,1,1)^{T}$ with $v_{0}=0.01/t_{f}$ for $0\leq t\leq t_{f}$. Remaining control parameters are such that $\Delta_{ii}(t)=0$ for $i\geq 2$ and $\Omega_{ij}^{I}(t)=0$ at all times. Other numerical parameters and line conventions are identical to those of Fig. 1.
IV CONCLUSION
@@
To conclude, we have adapted the derivation of the quantum speed limit developed by Vaidmann [3] for the evolution of single states in the context of quantum unitary operators. This derivation highlights that the quantum operators must satisfy specific geometric conditions. In particular, the time derivative of the evolution operator must always lie in a plane spanned by the initial operator and the current operator in order to saturate the MT-QSL. Since the evolution operator must also remain unitary at all times, this imposes a stringent constraint, and the MT-QSL cannot always be saturated. We have developed an optimal control algorithm to design drivings that take into account geometric constraints to approach at best the QSL. More generally, the distance of the evolution operator (or the instantaneous quantum state) from this plane provides a quantitative measure of the deviation from QSL optimality. Our method could be effectively applied to improve the QSL efficiency in a wide variety of shortcut-to-adiabaticity techniques [40]. Since the existence of a speed limit is fundamentally connected to energy and entropy exchange rates through a time-information uncertainty relation [41], this methodology extends beyond quantum systems to include, for instance, classical stochastic systems governed by Liouvillian dynamics [17, 18].
@@
Acknowledgements We thank the support from the ANR project QuCoBEC (ANR-22-CE47-0008-02) and from the CAPES-COFECUB (20232475706P) program. F.I. acknowledges support from the Brazilian agencies CNPq (305638/2023-8), FAPERJ (210.570/2024), and INCT-IQ (465469/2014-0).
APPENDIX: DERIVATION OF THE EQUATION FOR THE ADJOINT OPERATOR DYNAMICS
@@
Let us now put the dynamical equation for the adjoint $\mathbf{p}(t)$ in a more explicit form. From Eqs.(31,32), one finds:
@@
$\>{}^{T}\mathbf{P}(t)\left(\begin{array}[]{cc}H_{I}(\mathbf{v}(t))&H_{R}(\mathbf{v}(t))\\ -H_{R}(\mathbf{v}(t))&H_{I}(\mathbf{v}(t))\end{array}\right)\mathbf{X}(t).$ (42)
@@
We have used the Hermiticity of the Hamiltonian, i.e. $H_{R}(\mathbf{v}(t))^{T}=H_{R}(\mathbf{v}(t))$ and $H_{I}(\mathbf{v}(t))^{T}=-H_{I}(\mathbf{v}(t))$. The second term in the dynamical equation for the adjoint vector $\mathbf{P}(t)$ is identical to the term driving the operator $\mathbf{X}(t)$, and thus corresponds to the Hamiltonian evolution for the adjoint evolution operator $U^{P}(t)=U_{R}^{P}(t)+iU_{I}^{P}(t)$ constructed from the adjoint vector (setting $\mathbf{P}(t)=(U_{R}^{P}(t),U_{I}^{P}(t))^{T}$). Summing up, one arrives at Eq. (35), with the extra driving terms associated to the derivatives of the cost function listed below (the cost associated to the Hamiltonian variance does not contribute to the dynamics of the adjoint operator):
$\frac{\partial f_{\rm Plane}(\mathbf{U}(t),\mathbf{v}(t))}{\partial U_{R}}$ $=$ $-\frac{2}{d^{2}\Delta H^{2}}\left[\left({\rm Tr}[H_{R}U_{R}-H_{I}U_{I}]-\frac{1}{d}{\rm Tr}[H_{R}]{\rm Tr}[U_{R}]\right)\left(H_{R}-\frac{1}{d}{\rm Tr}[H_{R}]\mathbf{1}_{d}\right)\right.$
$\left.-\left({\rm Tr}[H_{R}U_{I}+H_{I}U_{R}]-\frac{1}{d}{\rm Tr}[H_{R}]{\rm Tr}[U_{I}]\right)H_{I}\right]-\frac{2}{d^{2}}{\rm Tr}[U_{R}]\mathbf{1}_{d},$
$\frac{\partial f_{\rm Plane}(\mathbf{U}(t),\mathbf{v}(t))}{\partial U_{I}}$ $=$ $-\frac{2}{d^{2}\Delta H^{2}}\left[\left({\rm Tr}[H_{R}U_{R}-H_{I}U_{I}]-\frac{1}{d}{\rm Tr}[H_{R}]{\rm Tr}[U_{R}]\right)H_{I}\right.$
$+$ $\left.\left({\rm Tr}[H_{R}U_{I}+H_{I}U_{R}]-\frac{1}{d}{\rm Tr}[H_{R}]{\rm Tr}[U_{I}]\right)\left(H_{R}-\frac{1}{d}{\rm Tr}[H_{R}]\mathbf{1}_{d}\right)\right]-\frac{2}{d^{2}}{\rm Tr}[U_{I}]\mathbf{1}_{d},$
$\frac{\partial f_{\rm Phase}(\mathbf{U}(t),\mathbf{v}(t))}{\partial U_{R}}$ $=$ $\frac{2}{d^{2}}\sqrt{f_{\rm Phase}(\mathbf{U}(t),\mathbf{v}(t))}\left(-{\rm Tr}[U_{I}]H_{I}+{\rm Tr}[U_{R}]H_{R}+{\rm Tr}[H_{R}U_{R}-H_{I}U_{I}]\mathbf{1}_{d}-\frac{2}{d}{\rm Tr}[U_{R}]{\rm Tr}[H_{R}]\right),$
$\frac{\partial f_{\rm Phase}(\mathbf{U}(t),\mathbf{v}(t))}{\partial U_{I}}$ $=$ $\frac{2}{d^{2}}\sqrt{f_{\rm Phase}(\mathbf{U}(t),\mathbf{v}(t))}\left({\rm Tr}[U_{I}]H_{R}+{\rm Tr}[U_{R}]H_{I}+{\rm Tr}[H_{R}U_{I}+H_{I}U_{R}]\mathbf{1}_{d}-\frac{2}{d}{\rm Tr}[U_{I}]{\rm Tr}[H_{R}]\right).$ (49)

### v2

@@ IV.1 Optimization in a 4-dimensional space with the Pontryagin Maximum Principle
These results demonstrate that achieving — or even approaching — the quantum speed limit critically depends on both the target gates and the available dynamical couplings. The procedure’s performance, specifically the trade-off between quantum speed and quantum fidelity, improves with the number of available control parameters, as the geometric optimization routine can then explore a broader range of quantum paths.
Gate Fidelity $J_{\rm Plane}[U,H]$ $J_{\rm Phase}[U,H]$ $\overline{\eta}$
$U_{\rm QFT}$ 0.9995 0.426 0.104 0.700
$U_{\rm``QFT^{\prime\prime}}$ 1.000 0.063 0.079 0.928
$U_{\rm Hadamard}$ 0.9998 0.507 0.0001 0.672
$U_{\rm CNOT}$ 0.654 0.066 0.063 0.935
Table 5: Implementation of quantum gates using fidelity as the sole optimization criterion ($p_{1}=p_{2}=0$) with a control limited to 7 couplings $\mathbf{v}(t)=(\Delta_{11}(t),\Omega_{ij}^{R}(t))$. Other numerical parameters: $100$ iterations, variable step (initially $\epsilon=0.25/t_{f}$, increased by $50\%$ every 15 iterations, decreased by $10\%$ in case of cost increase), with $N_{t}=1000$ time bins.
Gate Fidelity $J_{\rm Plane}[U,H]$ $J_{\rm Phase}[U,H]$ $\overline{\eta}$
$U_{\rm QFT}$ 0.953 0.0046 0.0035 0.960
$U_{\rm``QFT^{\prime\prime}}$ 0.9993 0.034 0.035 0.965
$U_{\rm Hadamard}$ 0.9762 0.349 0.041 0.744
$U_{\rm CNOT}$ 0.9974 0.281 0.011 0.806
Table 6: Fidelity optimization while minimizing both the plane and phase cost functionals with the same weight ($p_{1}=p_{2}=0.5$) with a control limited to 7 couplings $\mathbf{v}(t)=(\Delta_{11}(t),\Omega_{ij}^{R}(t))$. Up to $200$ iterations with a variable step as in Table 2 and $N_{t}=1000$ time bins.
IV.2 Geometric Optimization with the Chopped Random Basis quantum optimization
We now compare our geometric approach with CRAB (Chopped Randomized Basis) optimization. Differences in results are expected, as various optimization procedures can converge to different local minima. We again consider a time-dependent, $N_{C}$-dimensional control parameter $\mathbf{v}(t)$ representing the Hamiltonian couplings. Here, $N_{C}=N^{2}$ in an $N$-dimensional Hilbert space when all couplings can be independently controlled.
The core idea of the CRAB method is to approximate each time-dependent component $v_{i}(t)$ using an ansatz built from a truncated randomized basis. The most common example, and the one we use below, involves a finite set of Fourier modes with randomly chosen frequencies and phases. For each control parameter $v_{i}(t)$, we employ the following ansatz:
$v_{i}(t)=A_{i,0}+\sum_{n=1}^{N_{\omega}}\left[A_{i,n}\cos(\omega_{i,n}t)+B_{i,n}\sin(\omega_{i,n}t)\right].$ (54)
We’ve chosen a distinct set of frequencies $\{\omega_{1},...,\omega_{N_{\omega}}\}$ for each component $v_{i}(t)$, defined as $\omega_{i,n}=n\pi/t_{f}+\delta\omega_{i,n}$, where $\delta\omega_{i,n}=2\pi\delta_{i,n}/t_{f}$ (with $\delta_{i,n}\in[0,1]$). Randomization is a key ingredient for accelerating convergence. In our CRAB algorithm implementation, these frequencies are fixed during the initialization stage and then used throughout successive iterations.
Within this approach, each coupling $v_{i}(t)$ is represented by a finite-dimensional vector $(A_{i,0},A_{i,1},...,A_{i,N_{\omega}},B_{i,1},...,B_{i,N_{\omega}})$, as the set of frequencies $\omega_{i,n}$ is treated as fixed variables. The global cost functional $J[U,H]$ from Eq. (32) can therefore be viewed as a simpler function $\mathcal{F}(A_{i,0},A_{i,1},...,A_{i,N_{\omega}},B_{i,1},...,B_{i,N_{\omega}})$ of $N_{\rm tot}=(2N_{\omega}+1)N_{C}$ variables. Finding the optimal control path then reduces to a simpler problem of multi-variable function minimization. Increasing the number of modes in the randomized basis expands the space of functions explored by the algorithm, leading to a better approximation of the optimal solution, but also comes with significant computational overhead.
We present the results from the CRAB optimization using the cost functional with weights $(p_{1}=0.5,p_{2}=0.1)$, which yielded the best outcomes for this routine. The optimization was conducted using a direct-search method to minimize the multi-variable function. This function is derived by solving the Schrödinger equation for the evolution operator, utilizing the set of parameters $(A_{i,0},A_{i,1},...,A_{i,N_{\omega}},B_{i,1},...,B_{i,N_{\omega}})$ that define each coupling’s temporal profile.
In our case, we set $N_{\omega}=10$ and $N_{C}=16$, meaning the minimization was performed with respect to $N_{\rm tot}=336$ variables. As this is a computationally demanding task, we used a smaller number of time bins ($N_{t}=100$) than in the PMP optimization. Once the pulses were obtained from the optimization routine, we performed a linear interpolation and resolved the Schrödinger equation with $N_{t}=10^{4}$ time bins to ensure accurate estimates of the QSL and quantum fidelity.
The optimization results for the four quantum gates, $U_{\rm QFT}$, $U_{\rm``QFT^{\prime\prime}}$, $U_{\rm Hadamard}$ and $U_{\rm CNOT}$ are presented in Table 7. Each optimization consumed a total computing time of approximately 12 hours on a personal laptop, significantly more than the 20 minutes required for the Pontryagin optimization routine.
The trade-off between quantum speed limit and quantum fidelity from the CRAB optimization, as shown in Table 7, is slightly degraded compared to the Pontryagin optimization (see Table 2 and Table 4). We attribute this reduced performance to the slower convergence of the CRAB process when handling a large number of couplings. The results depend again on the considered target gate: the CRAB optimization yields an implementation close to the QSL saturation limit for the Hadamard gate, but only delivers a quantum path with a low QSL efficiency for the CNOT gate.
Gate Fidelity $J_{\rm Plane}[U,H]$ $J_{\rm Phase}[U,H]$ $\overline{\eta}$
$U_{\rm QFT}$ 0.9993 0.054 0.045 0.951
$U_{\rm``QFT^{\prime\prime}}$ 0.9992 0.051 0.054 0.952
$U_{\rm Hadamard}$ 0.9999 0.003 0.000 0.998
$U_{\rm CNOT}$ 0.9975 0.279 0.226 0.707
Table 7: Fidelity optimization with the CRAB routine while minimizing both QSL cost functionals with weights $(p_{1}=0.5,p_{2}=0.1)$ in a full control scenario over $16$ couplings in a 4-dimensional system. Optimization realized with a discretization over $N_{t}=100$ time bins, with $500.000$ iterations for the gates $U_{\rm QFT}$$U_{\rm``QFT^{\prime\prime}}$ and $700.000$ iterations for the gates $U_{\rm Hadamard}$$U_{\rm CNOT}$.
Figure 2: Implementation of the gate $U_{\rm``QFT^{\prime\prime}}$ after geometric CRAB optimization with a full Hamiltonian control $(p_{1}=0.5,p_{2}=0.1)$, 2nd line of Table 7]. (a):Instantaneous QSL efficiency $\eta(t)$ as a function of the rescaled time $t/t_{f}$. (b):Fidelity $\mathcal{F}=|\langle U_{\rm target},U(t)\rangle|^{2},$ running cost functions $f_{\rm Plane}(U(t),H(t))$ (dash-dotted blue line) and $f_{\rm Phase}(U(t),H(t))$ (dotted red line) as a function of the rescaled time $t/t_{f}$. (c):Dynamical couplings $\Omega_{12}^{R}(t)$ (dash-dotted blue line), $\Omega_{12}^{I}(t)$ (dotted blue line), $\Omega_{13}^{R}(t)$ (dash-dotted red line), $\Omega_{13}^{I}(t)$ (dotted red line), $\Omega_{23}^{R}(t)$ (dash-dotted green line), $\Omega_{23}^{I}(t)$ (dotted green line). (d):Dynamical couplings $\Omega_{23}^{R}(t)$ (dash-dotted blue line), $\Omega_{23}^{I}(t)$ (dotted blue line), $\Omega_{24}^{R}(t)$ (dash-dotted red line), $\Omega_{24}^{I}(t)$ (dotted red line), $\Omega_{34}^{R}(t)$ (dash-dotted green line), $\Omega_{34}^{I}(t)$ (dotted green line). (e): Dynamical couplings $\Delta_{11}(t)$ (dash-dotted blue line), $\Delta_{22}(t)$ (dash-dotted red line), $\Delta_{33}(t)$ (dotted green line), $\Delta_{44}(t)$ (dotted purple line). All couplings in (c,d,e) are given in units of $1/t_{f}$.(f):Matrix elements of the operator $\tilde{U}(t_{f})=U(t_{f})e^{-i\varphi}$ (Real and Imaginary parts), with $\varphi={\rm Arg}[\langle U_{\rm target},U(t_{f})\rangle]$.
V CONCLUSION
@@
To conclude, we have adapted the derivation of the quantum speed limit developed by Vaidmann [3] for the evolution of single states in the context of quantum unitary operators. This derivation highlights that the quantum operators must satisfy specific geometric conditions. Precisely, the time derivative of the evolution operator must always lie in a plane spanned by the initial operator and the current operator in order to saturate the MT-QSL, and simultaneously fulfill a phase condition. Since the evolution operator must also remain unitary at all times, this imposes stringent constraints, and the MT-QSL cannot always be saturated. More generally, the distance of the evolution operator (or the instantaneous quantum state) from this plane as well as the integrated phase provide a quantitative measure of the deviation from QSL optimality.
Building on these geometric considerations, we’ve developed an optimal control algorithm designed to account for geometric constraints to best approach the Quantum Speed Limit (QSL). We’ve tested our approach using two distinct optimization procedures: the Pontryagin Maximum Principle (PMP) and the Chopped Randomized Basis (CRAB) quantum optimization. We investigated various quantum gates relevant to quantum information processing in a two-qubit system.
Our optimization method significantly reduces the departure from QSL optimality while maintaining an excellent level of quantum fidelity. Typically, for the quantum gates we considered, a few parts per thousand in quantum fidelity can be traded for a few percent increase in quantum speed efficiency. Such QSL optimization could lead to significant improvements in quantum devices where rapid implementation of quantum gates is critical.
In quantum spaces of arbitrary dimensions, proving the existence of a quantum trajectory that reaches a given target quantum gate and saturates the QSL can be a challenging problem. By presenting suitable time-dependent quantum protocols that achieve the target gate within a close distance, our optimization procedure provides lower bounds on the optimal average QSL efficiency that can be attained.
Our method could be effectively applied in conjunction with shortcut-to-adiabaticity techniques to improve the QSL efficiency of such quantum protocols [47]. Given that the existence of a speed limit is fundamentally connected to energy and entropy exchange rates through a time-information uncertainty relation [48], this methodology extends beyond quantum systems. It can, for instance, include classical stochastic systems governed by Liouvillian dynamics [18, 19].
@@
Acknowledgements We thank the support from the Institut Universitaire de France, from the ANR project QuCoBEC (ANR-22-CE47-0008-02) and from the CAPES-COFECUB (20232475706P) program. F.I. acknowledges support from the Brazilian agencies CNPq (305638/2023-8), FAPERJ (210.570/2024), and INCT-IQ (465469/2014-0).
APPENDIX A: EXPRESSION OF THE COST FUNCTIONALS
The running plane and phase cost functions $f_{\rm Plane}(U(t),H(t),t)$ and $f_{\rm Phase}(U(t),H(t),t)$ [Equations (30) and (31)] are expressed below in terms of real-valued quantities. Let us define $Z(t)=\langle U(t),\mathbb{I}\rangle\langle\mathbb{I},U_{\perp}(t)\rangle$.
[…]
$-\frac{1}{d^{2}}\left({\rm Tr}[U_{R}]^{2}+{\rm Tr}[U_{I}]^{2}\right)$
${\rm Re}[Z(t)]$ $=$ $\frac{1}{d^{2}\Delta H}\left[{\rm Tr}[U_{I}]{\rm Tr}[H_{R}U_{I}+H_{I}U_{R}]+{\rm Tr}[U_{R}]{\rm Tr}[H_{R}U_{R}-H_{I}U_{I}]-\frac{1}{d}{\rm Tr}[H_{R}]\left({\rm Tr}[U_{R}]^{2}+{\rm Tr}[U_{I}]^{2}\right)\right].$
${\rm Im}[Z(t)]$ $=$ $\frac{1}{d^{2}\Delta H}\left[-{\rm Tr}[U_{I}]{\rm Tr}[H_{R}U_{R}-H_{I}U_{I}]+{\rm Tr}[U_{R}]{\rm Tr}[H_{I}U_{R}+H_{R}U_{I}]\right]$ (55)
With $|\langle\mathbb{I},U(t)\rangle|^{2}=\frac{1}{d^{2}}\left({\rm Tr}[U_{R}]^{2}+{\rm Tr}[U_{I}]^{2}\right)$, one readily obtains $f_{\rm Plane}(U(t),H(t))$ (30). The phase running cost function reads $f_{\rm Phase}(U(t),H(t))={\rm Re}[Z(t)]^{2}/({\rm Re}[Z(t)]^{2}+{\rm Im}[Z(t)]^{2})$. To compute the extra driving terms associated to the derivatives of the cost functional, we use that $\frac{d{\rm Tr}[AX]}{dX}=A^{T}$ (where T is the matrix transpose), together with the properties that $H_{R}^{T}=H_{R}$ and $H_{I}^{T}=-H_{I}$.
APPENDIX B: DERIVATION OF THE EQUATION FOR THE ADJOINT OPERATOR DYNAMICS
@@
Let us now put the dynamical equation for the adjoint $\mathbf{p}(t)$ in a more explicit form. From Eqs.(36,37), one finds:
@@
$\mathbf{P}(t)^{T}\left(\begin{array}[]{cc}H_{I}(\mathbf{v}(t))&H_{R}(\mathbf{v}(t))\\ -H_{R}(\mathbf{v}(t))&H_{I}(\mathbf{v}(t))\end{array}\right)\mathbf{X}(t).$ (59)
where $\mathbf{P}(t)^{T}$ is the transpose of $\mathbf{P}(t)$.
@@
We have used the Hermiticity of the Hamiltonian, i.e. $H_{R}(\mathbf{v}(t))^{T}=H_{R}(\mathbf{v}(t))$ and $H_{I}(\mathbf{v}(t))^{T}=-H_{I}(\mathbf{v}(t))$. The second term in the dynamical equation for the adjoint vector $\mathbf{P}(t)$ is identical to the term driving the operator $\mathbf{X}(t)$, and thus corresponds to the Hamiltonian evolution for the adjoint evolution operator $U^{P}(t)=U_{R}^{P}(t)+iU_{I}^{P}(t)$ constructed from the adjoint vector (setting $\mathbf{P}(t)=(U_{R}^{P}(t),U_{I}^{P}(t))^{T}$). Summing up, one arrives at Eq. (40).